A random sample of ?=80 in-coming college freshmen showed that 44 were planning to bring a car with them to campus next fall. Suppose that we are interested in forming a confidence interval at 95 percent for the proportion of all in-coming college freshmen who are planning to bring a car to campus next fall.

Where appropriate, express your answer as a proportion (not a percentage). Round answers to three decimal places.

Estimate: _______

Standard error: ________

Multiplier:_________
Answer #1

given data are:-

sample size(n) = 80 , x = 44

a).Estimate be :-

\hat{p}=\frac{x}{n}=\frac{44}{80}\approx \mathbf{0.55}

b). The Standard error be:-

SE=\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}=\sqrt{\frac{0.55(1-0.55)}{80}}\approx 0.056

c). the multiplier (z critical value for 95 % confidence level) be:-

z^*=1.96

the 95% confidence interval for the proportion of all in-coming college freshmen who are planning to bring a car to campus next fall be:-

=\hat{p}\pm z^**SE

=0.55\pm (1.96*0.056)

\approx \mathbf{(0.440,0.660)}

*** if you have any doubts about the problem please mention it in the comment box... if satisfied please LIKE.